### v1

# Role of symmetry in the correlated orientational disorder phase of entropic crystals

Abstract
Correlated orientational disordered phase is a thermodynamic phase with the constituent particles following some conserved laws or correlation despite the orientational disorder present in the system. Polyhedral particles in the presence of hard interaction serve the purpose of studying the orientational disorder in entropy driven self-assembly. We report the computational evidence of the role of symmetry of both particle and self-assembled crystal in the single-component system of anisotropic polyhedra in the correlated orientational disordered phase. The point groups of the particle and crystal structure appear to make a direct correspondence with the existence of multiple unique orientations of the system satisfying all the orientational conserved quantities towards the correlation. As a signature of correlation, the system exhibits orientationally frozen state at sufficiently higher packing fractions and discrete orientational hopping of the particles are observed at comparatively lower density solids. This symmetry relationship appears to control all the aspects of orientational correlation in the crystalline solids which might be useful to draw fundamental insights about the disordered crystalline solids with orientational correlation as well as designing the disorder in the crystals.
pacs:
Valid PACS appear hereI Introduction
Ordering of classical many-body systems is a spontaneous process in nature which occur in different length scales Whitesides and Boncheva (2002). Crystallization drives a disordered system to a translationally ordered state where all the constituents form a translational broken symmetry state as observed in the domain of colloidal crystals Pawel (1983); Dinsmore et al. (1998); Li et al. (2016); Orr et al. (2022); Boles et al. (2016), molecular crystals Timmermans (1961); Guihrie and Mccullough (1961); Reynolds (1975); Harfenist et al. (1996); Hochli et al. (1990); Nitta (1959) or nanoparticle superlattices Mirkin et al. (1996); Macfarlane et al. (1996); Whitesides and Boncheva (2002); Maye et al. (2007); Henzie et al. (2012); Bodnarchuk et al. (2011); Boneschanscher et al. (2014); Lin et al. (2017); Zhou et al. (2022); Geuchies et al. (2016); Shevchenko et al. (2006); Zhang et al. (2013); Jones et al. (2010); Glotzer and Solomon (2007). Crystal structure of rigid anisotropic particles exhibit different kinds of orientational phases due to the rotational motion of the particles while maintaining proper translational order Mirkin et al. (1996); Glotzer and Solomon (2007); O’Brien et al. (2016); Damasceno et al. (2012a); Gantapara et al. (2013). The possible orientational phases in the crystalline assemblies include the plastic crystals Timmermans (1961); Reynolds (1975); Lee et al. (2023); Karas et al. (2019); Dullens and Petukhov (2007); Meijer et al. (2017), orientationally ordered crystals Damasceno et al. (2012a); Karas et al. (2019); Gantapara et al. (2013); Lu et al. (2021); Henzie et al. (2012), correlated disordered crystals Keen and Goodwin (2015); Chaney et al. (2021); Meekel and Goodwin (2021) or orientational glass Loidl (1989); Hochli et al. (1990); Karas et al. (2019); Abbas et al. (2022); Vdovichenko et al. (2015). The existence of these orientational phases were reported at different length scales either in experiments Abbas et al. (2022); O’Brien et al. (2016); Jones et al. (2010) or computer simulations Damasceno et al. (2012a); Gantapara et al. (2015); John et al. (2008); Karas et al. (2019). Rotator or plastic crystal is a orientational disorder phase with a long-range translational order and very short-range orientational order Timmermans (1961); Reynolds (1975). All the particles can rotate freely in the orientational space while staying at the lattice sites of the crystal; such states can be termed as “freely rotating plastic crystals”. This phenomenon is well documented in the field of colloidal crystals Pawel (1983); Orr et al. (2022); Boles et al. (2016), molecular crystals Vdovichenko et al. (2015); Even et al. (2016); Beake et al. (2017); Hochli et al. (1990) or nanocrystal superlattices Shevchenko et al. (2006); Zhang et al. (2013); Damasceno et al. (2012b); Maye et al. (2007); Lee et al. (2023); Damasceno et al. (2012a). The existence of orientationally ordered crystal structure with the long-range positional and orientational order was also established in earlier literature Damasceno et al. (2012a); Gantapara et al. (2013); Lu et al. (2021); Henzie et al. (2012). Direct experimental observations of orientational order and disorder have been investigated in colloidal crystals Dullens and Petukhov (2007); Meijer et al. (2017) and nanocrystal superlattices Deng et al. (2020); Elbert et al. (2021); Abbas et al. (2022) very recently and studying the orientations of the particles in the crystal structures became one of promising fields in the contemporary research. Correlated orientationally disordered phase is another kind of orientational phase which is less explored compared to the former two phases Keen and Goodwin (2015); Simonov and Goodwin (2020); Kundu et al. (2023). The existence of the orientational correlation in the disordered crystalline solids was already reported in the molecular crystals Keen and Goodwin (2015) and became an intriguing field to realize the orientational correlation coupled with the translational order.
Entropy driven ordering became one of the promising fields in studying crystallization where the constituents interact via hard-core interaction only Onsager (1949); Alder and Wainwright (1957); Berryman (1983); Frenkel and Ladd (1984); Frenkel and Mulder (1985); Veerman and Frenkel (1992); McGrother et al. (1996); Bolhuis and Frenkel (1997). To realize the direct experimental observations of various orientational phases in the crystalline solids, hard anisotropic particles were used as a promising model of colloidal particles or nanocrystals. Assembly of hard anisotropic particles turned into scientific interest for studying the complex behavior of the building blocks or entropy driven crystallization. A lot of computational studies have shown the evidences of the assembly of hard particles into different kinds of crystalline structures Alder and Wainwright (1957); Bolhuis and Frenkel (1997); John et al. (2008); Haji-Akbari et al. (2009); Torquato and Jiao (2009); Damasceno et al. (2012a); Abbas et al. (2022). In particular, theoretical studies and computer simulations on freezing and melting of hard spheres, sphero-cylinders etc. laid the foundation to study crystallization, where entropy is considered as only driving force Alder and Wainwright (1957); Frenkel and Ladd (1984); Frenkel and Mulder (1985); Veerman and Frenkel (1992). Exhaustive simulations of hard convex polyhedra produced simple and complex crystal structures Damasceno et al. (2012a); Agarwal and Escobedo (2011); Gantapara et al. (2013, 2015); John et al. (2008); Ni et al. (2012) showing similarity in the assembled states achieved in the experiments of colloids and nanoscience Alloyeau et al. (2009); Abbas et al. (2022); Gong et al. (2017); Even et al. (2016); Ong et al. (2017); Meijer et al. (2017). Apart from studying the coupling between the translational and rotational motions of the particles, different kinds of orientational phases were also investigated using the anisotropic polyhedra as the idealized models of entropy driven assembly in both two and three dimensions Gantapara et al. (2015); Shen et al. (2019); Karas et al. (2019); Damasceno et al. (2012a); Lee et al. (2023); Karas et al. (2019). Studying the orientational disorder in the crystals of anisotropic polyhedra came under scrutiny to serve the purpose of understanding the disordered phases in a more deeper way. In a recent study, our investigation shed the light on the existence of the correlated orientational disordered phase in the crystalline solids, where the polyhedral particles were found to assemble into orientationally disordered crystal structures while maintaining the orientational correlation, completely driven by the entropy. The signature of correlations in the crystalline solid was reported but the inherent orientational disorder under the effect of correlation is still unreported in the simplest model system which demands further attention to understand the essence of the phase thoroughly.
Correlated orientational disordered phase was realized as a disordered crystal with multiple unique orientations occurring at the comparatively higher packing fractions and maintaining some nontrivial forms of the correlation Keen and Goodwin (2015); Simonov and Goodwin (2020); Kundu et al. (2023). Multiple unique orientations of the particles were already reported in the colloidal crystals or nanocrystal superlattices in both experiments Abbas et al. (2022) and computer simulations Shen et al. (2019); Karas et al. (2019); Lee et al. (2023) but lacked the insight of correlation. Our previous computational study reported the existence of few orientational attributes which were preserved in the disordered crystalline solids. We argued the collective occurrence of all orientational attributes as the signature of strong orientational correlation Kundu et al. (2023). The phase persisted within certain range of packing fractions, where the orientational motion of the particles were frozen at the densest states and discretely mobile within the unique orientations at comparatively lower packing fractions. The discrete hopping of anisotropic polyhedra within the unique orientations were observed multiple times in the crystalline solids of hard polyhedra in both two Shen et al. (2019) or three dimensions Lee et al. (2023). We also anticipated the “symmetry mismatch” between the non-crystallographic point group of the polyhedral particle and point group of the crystallographic symmetry responsible for the existence of the correlated disordered phase. This gave rise of an obvious question whether the point groups of the particle and crystal could be coupled in some forms to get a better handle of the orientational correlation under the translational broken symmetry state in three dimensions.
In this article, we present the computational study of the single-component systems of hard convex polyhedra to investigate the role of symmetry of the particle and crystal structure in the presence of orientational correlation. The chosen polyhedra were Elongated Pentagonal Dipyramid (EPD), Elongated Square Gyrobicupola (ESG) and Elongated Pentagonal Gyrocupolarotunda (EPG) which were reported to form face centred cubic (FCC) crystals with no positional transitions in the entire solid regions Damasceno et al. (2012a); Karas et al. (2019); Kundu et al. (2023). The plastic crystal phases appear for all the shapes at lower packing fractions before the transition to the isotropic liquid phase. The existence of the correlated disordered phase was reported for all the shapes in our previous study Kundu et al. (2023). In this research, we explore the interplay between the particle symmetry and crystallographic symmetry and observe the explicit role of both symmetries in controlling the correlated disordered phase by preserving all the orientational attributes as discussed in the previous study. Our findings might play an important role in developing fundamental theory to explain such broken symmetry states in the many-body systems or designing new materials with correlated disorder.
II Models and methods
We performed hard particle Monte Carlo simulations (HPMC) using HOOMD-Blue toolkit Anderson et al. (2016), where the metropolis method was determined by the overlapping of the particles . A dilute system was prepared followed by slow volume compression of the system upto 0.55 or 0.6, depending on the shapes. After the equilibration performed in constant volume simulation for sufficiently large Monte Carlo steps, the crystal structure was obtained. The pressure was computed in the constant volume simulation using the scale distribution function implemented in HOOMD. The system was compressed by increasing the pressure value slowly and allowed to equilibrate for 100 Monte Carlo (MC) steps under isobaric-isothermal ensemble (NPT). This protocol helped us to achieve the system at very high packing fraction after which no significant changes in system were observed. The entire compression simulation was performed at 25-30 pressure values consecutively where the system at highest achievable packing fraction was chosen for the melting simulations. We carried out the melting simulations under NPT ensemble at the same pressure values which were used during the compression runs. This was continued until the isotropic liquid phase was achieved. The simulation method granted the permission to analyze the crystal structures and the orientations of the particles at packing fractions during the compression or expansion of the system.
Single particle orientation is represented by quaternion in the system. The orientational difference between two particles is described as the minimum of all angles where and are the quaternions of particle and respectively. The proper point group of the particle () was considered while calculating the pairwise angles between any two particles. Two particles are said to be perfectly ordered if = 0. The distribution of all pairwise angles in the system was calculated measuring the overall angular dispositions of the particles but it failed to disclose the number of unique orientations in the system. To identify the unique orientations, a separate recipe was followed by choosing three arbitrary quaternions form the system , , maintaining a finite pairwise angle among themselves. For the -th particle chosen from the system, three angles with the references corresponded to a set consisting of the angles . In this way, the sets of three angles were calculated for all particles in the system exhibiting a three dimensional distribution of the angles. In presence of statistical noise, the distribution produced multiple clusters in the angular space where the total number of clusters indicated the number of unique orientations in the system. The particles belonging to a same cluster were considered as orientationally ordered and the particles from different clusters were accounted as orientationally disordered.
In order to analyze the orientations of the particles sitting at the lattice sites of the crystal, calculating the spatial arrangements of the orientations was necessary. To get a better handle, the rotational symmetry axes of both the particle point group and the crystallographic point group were considered to see the alignments. To accomplish the information, the unit cells of the crystal structure along with the particle orientations at each lattice sites were obtained. We defined the orientation of a unit cell by an orthogonal rotation matrix in the global frame followed by the detection of the set of rotational symmetry operations (both the rotational axes and angles) of the crystallographic point group in the local frame of unit cell. All rotational symmetry axes of the particle were obtained from the point group in the local frame. We focused on the highest order rotational symmetry axes of the particle and all rotational symmetry axes of unit cell of the crystal structure which were defined by the sets , respectively in the corresponding local frames. The highest order rotational symmetry axes of the particle explicitly meant that if the particle point group contains few proper rotational operations of , and , then the axes corresponding to the operation are said to be highest order rotational symmetry axes of the particle.
To establish any relationships among the rotational symmetry axes of the particle point group and crystallographic point group, the reference frame was required to be fixed for all the axes. Considering the global frame as reference, we carried out the orthogonal transformations from the local frame of particle and unit cell to the global frame using the Equations 1, 2;
where was the rotational matrix defined in the global frame corresponding to the unit quaternion of -th particle. and were one of the elements of the sets , respectively with [, ] and and [, ]. and corresponded to the total number of highest order rotational axes of the particles and all rotational symmetry axes of the unit cell. We defined and as two separate sets containing both kind of axes in the global reference frame. The set corresponding to all rotational symmetry axes of the crystal was categorized according to the particular kind of possible rotational operations in the crystallographic point group like , , etc. For each particle , the angles between one of the highest order rotational axes from the set and the rotational symmetry axes in the set for each kind of rotational operation were measured as , , etc. The alignment of any particle axes with the , , axes of crystal symmetry can also be estimated as , , respectively when all these operations exist in the crystallographic point group. Again, was calculated for -th particle as the minimum of all the angles formed between the -th element of the set and all rotational axes of the crystal irrespective of any kind of operation according to the Equation 3.
where and were the indices of highest order rotational axes of the particle
and all rotational axes of the crystal chosen from the sets and respectively. The minimum angle was calculated for all the particles in the system () which was defined as a set and corresponded to the total number of particles in the system. The data of was presented as a two-dimensional histogram to analyze the alignment of the particles in the unit cell of the crystal structure. As the definition of suggests, directional matching between the highest order particle symmetry axis and any particular kind of crystal symmetry axes (, , ) gets reflected in the distribution of . The distribution can be considered as an order parameter to distinguish any different kind of orientational behavior from the complete random disordered state i.e., plastic crystal phase. In principle, for a crystalline state with random orientations of the particles, the distribution is expected to be Gaussian in the equilibrium phase. Any significant deviation from the Gaussian distribution may serve the purpose of identifying other kind of orientational characteristics in the system.
III Results and Discussion
III.1 Correlated orientational disorder phase
The systems of Elongated Pentagonal Dipyramid (EPD), Elongated Square Gyrobicupola (ESG) and Elongated Pentagonal Gyrocupolarotunda (EPG) shapes were simulated where all the shapes were reported to self-assemble into orientational correlated disordered phases under FCC crystalline symmetry at sufficiently higher pressure regions. In the crystalline states, all the particles followed a few conserved orientational attributes as a signature of orientational correlation which was present in the system without any particular spatial arrangements. In the previous research, it was argued that the crystal structure appeared to be orientationally disordered in this phase due to the existence of a large number of FCC unit cells which were classified based on the orientations of the particles. A straightforward notion of orientational order and disorder was defined to analyze the orientations of anisotropic particles in the crystal structure and it was shown that the production of a complete crystal structure was impossible from the knowledge of a single unit cell. The disordered crystalline phase in the presence of orientational correlation was termed as “correlated orientational disordered” phase. Following the earlier research, the simulation snapshots of the EPD, ESG and EPG shapes at the highest achievable packing fractions in self-assembly are shown in the first row of Figs. 1A, B, C respectively. The existence of finite number of unique orientations in the systems are showcased as clusters in the three dimensional orientational space spanned by , and as depicted by the second row of Figs. 1A, B, C. The number of clusters i.e., unique orientations appeared to be four, six and eight for the systems of EPD, ESG and EPG shapes respectively. The distributions of all pairwise angles maintained fixed orientational differences for the systems and all the particles got equally compartmentalized within the unique orientations (third row of Fig. 1) as a signature of orientational correlation ruling out the simultaneous occurrence of all these possibilities as a coincidence. The particles were colored according to the unique orientations depicted by the clusters in three dimensional orientational space and the same snapshots shown in the first row of Fig. 1 are represented with colored particles in the last row of Fig. 1. Though the exact form of correlation function was unknown, the discrete orientational hopping of the particles at slightly lower packing fractions than the densest states was observed among the unique orientations with equal probability. These observations strengthened the existence of a strong orientational correlation in the crystalline states.
III.2 Alignment of the rotational symmetry axes of particle point group and crystallographic point group
To understand the orientational alignments of the particles sitting at the lattice sites of the FCC unit cell in a more precise way, the particle symmetry and the crystallographic symmetry were taken into account simultaneously. As the orientations of the particles in the correlated disordered phase appeared to be explicitly dependent on the crystal structure, the consideration of the crystallographic point group played an important role to realize the orientational correlation. We explored all kinds of operations in the point groups of both the particle and crystal structure but no other operations except for the proper rotations appeared to be relevant in the correlation. Among all the proper rotational operations, only the highest order rotational axes of the particle seemed to maintain a direct relationship with the rotational axes of the corresponding crystallographic symmetry. The finite set of unique orientations was perceived in the context of the alignment between the directions of the axes in the sets and . For each shape used in this study, there are only two rotational symmetry axes of the particle point group corresponding to the highest order if two opposite directions are considered as non-equivalent, i.e., axes for EPD (point group - ), axes for ESG (point group - ) and EPG (point group - ) shape shape. The EPD shape contains another 5 proper rotational axes and ESG shape contains , 4 rotational axes as the lower order symmetry in the corresponding point group, whereas no lower order rotational symmetry exists in the point group of EPG shape. As all three systems self-assembled into FCC crystals, the crystallographic point group appeared to be under the Cubic class consisting of the proper rotational operations 8, 6, 6, 3. We estimated the angles , , for all particles in the system to uncover the alignments of the rotational symmetry axes of particle point group and crystallographic point group.
In the three dimensional space, a vector indicates a unique direction and the alignment can be realized by the intermediate angle between these two vectors. If the intermediate angle between the two directions is zero then the vectors are exactly parallel; any finite value of the angle suggests the misalignment of the two vectors. This idea was used to decipher the alignments of the particles sitting at the lattice sites of the FCC unit cell upon successful identification of the unit cells of the systems. The directions of the rotational symmetry axes of the particles were measured with the directions of the rotational symmetry axes of the corresponding crystal structure to recognize the alignments of the particles in the unit cell. The alignments were realized by the ensemble averaged distributions of , and calculated between the highest order particle rotational symmetry axes and crystallographic axes, as the symmetry contains only the , and rotational operations. In the distribution of , all the axes were considered without differentiating the axes from , where both the operations were present in point group. As both the EPD and EPG shapes contain axes as the highest order, the distributions of angles calculated between the axes of the particle point group and only axes of the crystallographic point group, exhibit the peaks around indicating the alignments to be parallel barring the statistical noise as shown in the second row of Figs. 2A, C. The other distributions for these shapes, and display the peaks at and estimated between the axes of the particles and all the and axes of the unit cell respectively ruling out the possibility of these crystal symmetry axes in the alignment of the particles. Similar kind of analysis shown in Fig. 2B suggests the alignments of the highest order axes of ESG shapes to be parallel with both the and axes of the FCC unit cell exhibiting the angles at for both cases. The distribution of produced a peak at suggesting the misalignment of the axes of the ESG shape with all axes of crystal. In the correlated disordered phase, any axis of EPD or EPG shape was observed to be aligned with the direction of any axis of the FCC unit cell consisting of total eight axes. These analysis followed that total eight unique directions were expected to occur for each of the EPD and ESG systems in the correlated disordered phase where the other lower order rotational symmetry axes of the particles failed to produce any alignments with the crystallographic rotational symmetry axes, if existed (see SI Fig. 11 for details). But in the systems of EPD and EPG shapes, the number of unique orientations appeared to be four and eight respectively as showcased in the second row of Figs. 1A, C which will be discussed later. Similarly, the directions of axes of ESG shape appeared to be aligned with any or axes of the crystallographic point group. As the point group contains 6, 3, 6 as proper rotational operations, total eighteen directions of particle axes were possible in the ESG system considering two exactly opposite axes as distinguishable. The appearance of only six unique orientations in the ESG system as observed in the Fig. 1B will be justified later in the context of the directions of highest order particle axes. Similarly, the other lower order axes of ESG shapes were not found to play any role in the alignment of the particles as found for the EPD shapes (SI Fig. 11).
The ensemble averaged distributions of calculated between the axes of the sets and were observed at the densest states corresponding to the correlated disordered phase as histograms in two dimensions. Considering all the crystallographic axes of point group as distinguishable, the distribution of appeared as an alternate representation of the plots shown in Fig. 2 with a major difference in identifying the responsible crystallographic axes in the alignments of the particles. In the distributions shown in Fig. 3A, B, C, all the crystallographic symmetry axes () were categorized so that (0, 5), (6, 17) and (18, 25) corresponded to , and axes of point group respectively, where the three axes were not shown separately as those coincided with the axes. For EPD and EPG shapes, the axes of the crystal symmetry were responsible in the alignment of the highest order particle symmetry axes forming the angles within as indicated in Figs. 3A, C. The color bar represents the normalized population density of particles. It was noticeable that all the eight crystallographic axes were equally populated by the particles in the system barring the statistical noise and showing the negligible effect of any other axes in the crystal. As the Fig. 3B suggests, the highest order particle symmetry axes of ESG shape appeared to be aligned with the six crystallographic axes where all the particles were also equally compartmentalized within the six directions satisfying the appearance of six unique orientations in the ESG system. This analysis suggested that the appearance of the peak near in the distribution of occurred as shown in Fig. 1B, because the three axes in the point group coincided with the axes. It was straightforward for the ESG shapes that if the axes of the particles aligned in the directions of axes of the FCC crystal, those axes of the particle were perpendicular to the planes of the crystal. Though this case happened for the ESG shapes, it did not turn out to be general for all the shapes exhibiting the correlated disordered phase like EPD and EPG systems. As the analysis suggested, the alignment of the rotational axes between the highest order particle symmetry and crystallographic symmetry turned out to be a characteristic feature of the phase in the presence of orientational correlation in the disordered crystalline structures. In the same way, the distributions of in the plastic crystal phases of the three shapes were observed which appeared to be almost Gaussian in nature showing distinct differences from the correlated disordered phases as shown in SI Fig. 7.
The visual representation of the alignment of the axes of EPD shapes in the unit cell of FCC crystals is shown in Fig. 4 at 0.74. In Fig. 4A, an ideal representation of the cube is shown with , , proper rotational symmetry axes and all mirror planes corresponding to the crystallographic point group symmetry. Fig. 4B suggests three randomly chosen corner particles and three face particles among fourteen particles of the FCC unit cell, identified randomly from the EPD system. The EPD shape contains point group and the axes of the EPD shapes aligned with the axes of the crystal as discussed earlier. Each polyhedral particle is shown along with all rotational symmetry axes and only two axes of the crystal are displayed in the unit cell which form or with the axes of the particles. The corresponding (hkl) planes responsible in the alignment of the particles in the unit cell are also displayed and marked as text. It could be realized that the directions of any rotational axes of point group were always normal to any of these planes shown in Fig. 4. The randomly chosen FCC unit cells for the three systems following the same color codes of the particles and all axes were shown in Fig. 5 suggesting the alignment of the particle axes in the unit cells of the crystal structures as discussed earlier. All the fourteen EPD particles located at the lattice sites of the same FCC unit cell are shown in SI Fig. 8 separately. The data of the other two FCC forming shapes ESG and EPG; are presented in the SI Figs. 9,10 in the similar fashion. Our observation regarding the straightforward relationships between the other rotational symmetry axes of the particle except the highest order and the crystallographic rotational symmetry axes in the context of unique orientations while maintaining the correlation was also confirmed visually by the Fig. 4.
To explore the orientationally ordered phase in the context of the symmetry of the particles and crystals, we performed the simulations of Truncated Cuboctahedron (TC) shape which was reported to form orientationally ordered body-centred tetragonal (BCT) structure with two unique orientations at the densest state Karas et al. (2019). It was known for the TC system consisting of two unique orientations with the orientational difference of 45∘. As TC shape contains point group symmetry, six axes appears to be the highest order rotational axes (considering two axes separated by 180∘ as non-equivalent). In the orientationally ordered TC system, no such crystallographic axis was found to be responsible in the parallel alignment of the particles defining the orientational order (see SI Fig. 12 for details). In the orientationally ordered crystal structure, the pairwise angle differences among the unique orientations are expected to maintain certain values depending on the crystal translational symmetry in case of the occurrence of multiple orientations, but any such relationship between the particle and crystallographic point group may not appear as a defining feature of the orientationally ordered phase as observed in the correlated disordered phase.
III.3 Discrete mobility of the highest order rotational axes of the particles
As the signature of correlation, the particles exhibited discrete orientational hopping within the unique orientations upon melting from the densest state, as reported in the previous study. Such kind of discrete motions were reported earlier in the systems of hard polygons where the phase was termed as “discrete plastic crystal” phase Shen et al. (2019). Our previous observation suggested, this behavior persisted before the arrival of the plastic crystal phase with complete freely rotating particles. One possible interpretation of the orientational hopping could be anticipated as the directional change of highest order rotational axis of the particles. As the orientational hopping always occurred within the unique orientations, it was expected that the highest order particle symmetry axes would toggle within a few fixed directions. We explored the discrete orientational hopping of the particles in the presence of orientational correlation in the light of highest order rotational axes of the particle point group. Considering the similar categorization of the axes as earlier, the discrete hopping of the axes were analyzed using the transition matrix averaged over all the particles and the trajectory for the three shapes as shown in Fig. 6. Each value in the two dimensional matrix indicated the hopping probability of axes within the corresponding crystallographic axes . The orientational motion of the particles were frozen at the highest packing fractions leading to the states where the axes were frozen giving rise to the population densities at the diagonal elements of the transition matrices only as shown in the first row of Figs. 6A, B, C. The axes of EPD, EPG shapes and axes of the ESG shapes aligned with and rotational axes of the crystal respectively. As the packing fraction decreased, the particles started to exercise the rotational motion by populating the off-diagonal elements of the transition matrices while the translational order of the system remained intact. The difference in the population densities between the diagonal and off-diagonal elements was noticed if the packing fraction was decreased slightly from the densest states as shown in second row of Figs. 6A, B, C. The differences in the ensemble averaged transition matrices suggested, maximum particles remained frozen at their respective orientations while a few particles changed the directions of axes by keeping the alignments fixed with any of the corresponding crystal axes. Upon further the melting of the systems, the differences minimized leading to the almost homogeneous distributions in the transition matrices which confirmed the changes of the directions occurred with equal probability for all particles in the system barring the statistical noise as shown in the last row of Figs. 6A, B, C for three shapes. Upon the melting of the systems, the analysis implied that the highest order particle rotational axes started to jump within the unique directions of the corresponding crystallographic axes discretely ruling out the possibility of any continuous change of the directions in random manner. For the ESG shapes, we failed to observe the two-step correlation in the transition matrices as discussed in the earlier research. The absence of the two-step correlation will be discussed later in support of our observations. The discrete directional hopping suggested that the changes of the axes always followed the correlation present in the system and did not exhibit any arbitrary behavior. This showed the existence of strong orientational correlation in the crystalline solids where all the orientational attributes were fixed as discussed in Fig. 1.
IV Conclusion
We report the computational evidence of the role of both particle point group and crystallographic point group in controlling the correlated orientational disorder phase in the entropic crystals. The disorder is maintained in the crystalline system without any spatial arrangements where a few orientational attributes are conserved following some complex forms of correlation. In this research, we discuss the orientational correlation from the symmetry aspect of both the particle and crystal structure. The appearance of the finite number of unique orientations in the system was realized in the context of particle and crystallographic point group. For EPD system, the number of unique orientations appeared to be four where as the axes of the EPD shapes were aligned in the directions of eight crystallographic axes. As the point group of EPD shape contains a plane, the shape is identical about the plane perpendicular to the axes of the shape. So, two axes separated by are considered as orientationally equivalent. As a result, two opposite directions of axes of the EPD shape separated by , are equivalent. Consequently, the number of unique orientations turns down to four from eight for EPD shape. This argument is straightforward and can be verified for the EPG shape containing point group with no plane leading to the eight unique orientations in the system. For the ESG shape with point group not containing any plane, the number of unique orientations appeared to be six, which was also consistent with the alignment of the axes of the particles with six axes of the crystal structure. The possible unique orientations of the shapes in the correlated disordered phase can be described by combining the point groups of both the particle and crystal structure. One obvious interpretation of the equal compartmentalization of the particles within the unique orientations is, the equal probability of the crystallographic symmetry axes in defining the alignments of the particles in the unit cells. As a particular kind of rotational symmetry axes of the crystal structure are separated by fixed angles, the directions of the highest order particle symmetry axes maintain certain angular differences in three dimensions leading to a fixed difference in the quaternions angles appearing in the distribution of all pairwise angles as shown in the third row of Fig. 1. All the conserved attributes corresponding to the correlated disordered phase can be explained in terms of the point group of both particle and crystal structure. For the shapes reported in this study, there exists only two directions of highest order rotational symmetry axes for the particle if the two opposite directions separated by are considered as non-equivalent i.e., axis for EPD and EPG shape and axis for ESG shape. The shapes having multiple highest order particle symmetry axes will be considered in the future investigations. Here, it turns out to be any particular kind of crystallographic axes (any of the , and for the cubic structure) responsible in the alignment of the particles. In general, the alignments will involve multiple kinds of crystallographic axes if large number of unique orientations occur in the system and the possibilities are open for further studies.
In the ESG system, there was a clear signature of second order correlation while investigating the discrete orientational hopping at lower packing fractions as reported in the previous study. We found no such signatures in the transition matrix of the highest order particle axes for the ESG shapes upon melting of the system as shown in Fig. 6B. The reason was uncovered which suggested that the presence of unique orientations separated by (second row of Fig. 1B) does not affect the direction of axis of a ESG particle. If the hopping between two unique orientations separated by occurs, no change in the direction of axis will be observed. If the transition matrix is constructed using the direction of highest order symmetry axis, there will be no signature of second order correlation as shown in the Fig. 6B. The second order of correlation depends on the shapes and does not appear to be a common attribute observed in the correlated disordered phase. Combining both the translational and rotational symmetry of the corresponding crystal structure with the particle point group symmetry may be enough to describe the correlated disordered phase completely as uncovered by the investigations.
In the correlated disordered phase, the orientations of the particles in the system, was realized in the context of the rotational symmetry of the particle and underlying crystal structure. A few orientational attributes were preserved as the effects of strong correlation which remained intact at the lower packing fractions also while the discrete orientational hopping were observed in the system. This phenomena was confirmed by the Fig. 6. Based on the different kinds of orientational behavior, the two phases can be distinguished. From our investigations, the following conclusions are inherited: (1) in the orientational ordered phase, highest order rotational symmetry axes of the particle point group form any finite angles with the rotational axes of the crystallographic point group, but certain pairwise angle differences among the the unique orientations were present if multiple unique orientations existed in the system, (2) the highest order rotational symmetry axes of the particles align along the directions of any rotational axes of the underlying crystal structure in the presence of orientational correlation in a disordered crystal, (3) the particles rotate freely maintaining no such orientational constraints with the underlying crystal structure in the plastic crystal phase. However, we did not observe the effect of other attributes except the particle and crystal symmetry responsible for the correlated disordered phase in the entropic crystal. The current investigation authorizes the combined role of both kind of symmetries up to a certain extent, which turns out to be a general characteristics exhibited by all particles in the system. According to our observations, single particle symmetry can be coupled with the crystalline symmetry to find the complex nature of the orientational correlation as reported in this study.
Our analyses reveal a new angle of studying the correlated disordered phase at the single particle level combining of the symmetry of the crystal structure. We established the role of translational and rotational symmetry of the crystal structure in controlling the disordered phase with orientational correlation in the entropy driven crystallization. A few similar kind of signatures were already mentioned in the molecular crystal literature where the phase was termed as “orientational glass” phase. The phase was reported to occur when particle symmetry did not match with the crystal symmetry and the essence of this phase was required to understand from the dynamical point of view. In the entropy driven assembly, we observe the existence of correlated disordered phase which satisfies every aspect and signatures of equilibrium phase. The dynamics of the particles was investigated in this study in the form of transition matrix which also suggested the signature of strong correlation. In this case, we believe that the dynamics became slower in this phase as the mobility of the particles slowly decreased as the packing fraction became higher. Beyond the plastic crystal phase, all the particles followed the orientational correlation with a strong relationship with the crystal structures and exhibited similar kind of qualitative behavior irrespective of the lower or higher packing fractions in this phase. All the signatures collectively make this a separate thermodynamic phase in the entropy driven assembly. As our observation suggests, this equilibrium phase consists of particles to preserve some orientational attributes across the system, where some complex form of correlations are present despite the slow dynamics of the system. Our research uncovered a straightforward relationship between the particle and crystal symmetry using one of the simplest model systems which might be useful in the context of designing principle of the correlated disordered phase in the material science or understanding the fundamental aspects of correlation present in the disordered systems in the presence of translational broken symmetry states.
Acknowledgements.
We wish to acknowledge financial support from DST-INSPIRE Fellowship provided to SK. AD thanks DST-SERB Ramanujan Fellowship (SB/S2/RJN-129/2016) and IACS start-up grant. KC thanks IACS for financial support. Computational resources were provided by IACS HPC cluster and partial use of equipment procured under SERB CORE Grant No. CRG/2019/006418.References
- Whitesides and Boncheva (2002) G. M. Whitesides and M. Boncheva, Proceedings of the National Academy of Sciences of the United States of America 99, 4769 (2002).
- Pawel (1983) P. Pawel, Contemporary Physics 24, 25 (1983).
- Dinsmore et al. (1998) A. D. Dinsmore, J. C. Crocker, and A. G. Yodh, Current Opinion in Colloid and Interface Science 3, 5 (1998).
- Li et al. (2016) B. Li, D. Zhou, and Y. Han, “Assembly and phase transitions of colloidal crystals,” (2016).
- Orr et al. (2022) N. H. Orr, T. Yanagishima, I. P. Dolbnya, A. V. Petukhov, and R. P. Dullens, Journal of Chemical Physics 157 (2022), 10.1063/5.0112602.
- Boles et al. (2016) M. A. Boles, M. Engel, and D. V. Talapin, Chemical Reviews 116, 11220 (2016).
- Timmermans (1961) J. Timmermans, J. Phys. Chem. Solids 18, 1 (1961).
- Guihrie and Mccullough (1961) G. B. Guihrie and J. P. Mccullough, J. Phys. Chem. Solids Pergamon Press 18, 53 (1961).
- Reynolds (1975) P. A. Reynolds, Molecular Physics 29, 519 (1975).
- Harfenist et al. (1996) S. A. Harfenist, Z. L. Wang, M. M. Alvarez, I. Vezmar, and R. L. Whetten, “Highly oriented molecular ag nanocrystal arrays,” (1996).
- Hochli et al. (1990) U. T. Hochli, K. Knorr, and A. Loidl, Advances in Physics 39, 405 (1990).
- Nitta (1959) I. Nitta, Zeitschrift fur Kristallographie, Bd. 112, 234 (1959).
- Mirkin et al. (1996) C. A. Mirkin, R. L. Letsinger, R. C. Mucic, and J. J. Storhoff, Nature 382, 607 (1996).
- Macfarlane et al. (1996) R. J. Macfarlane, B. Lee, M. R. Jones, N. Harris, G. C. Schatz, and C. A. Mirkin, Science 334, 204 (1996).
- Maye et al. (2007) M. M. Maye, D. Nykypanchuk, D. V. D. Lelie, and O. Gang, Small 3, 1678 (2007).
- Henzie et al. (2012) J. Henzie, M. Grunwald, A. Widmer-Cooper, P. L. Geissler, and P. Yang, Nature Materials 11, 131 (2012).
- Bodnarchuk et al. (2011) M. I. Bodnarchuk, E. V. Shevchenko, and D. V. Talapin, Journal of the American Chemical Society 133, 20837 (2011).
- Boneschanscher et al. (2014) M. P. Boneschanscher, W. H. Evers, J. J. Geuchies, T. Altantzis, B. Goris, F. T. Rabouw, S. A. V. Rossum, H. S. V. D. Zant, L. D. Siebbeles, G. V. Tendeloo, I. Swart, J. Hilhorst, A. V. Petukhov, S. Bals, and D. Vanmaekelbergh, Science 344, 1377 (2014).
- Lin et al. (2017) H. Lin, S. Lee, L. Sun, M. Spellings, M. Engel, S. C. Glotzer, and C. A. Mirkin, Science 355, 931 (2017).
- Zhou et al. (2022) S. Zhou, J. Li, J. Lu, H. Liu, J. Y. Kim, A. Kim, L. Yao, C. Liu, C. Qian, Z. D. Hood, X. Lin, W. Chen, T. E. Gage, I. Arslan, A. Travesset, K. Sun, N. A. Kotov, and Q. Chen, Nature 612, 259 (2022).
- Geuchies et al. (2016) J. J. Geuchies, C. V. Overbeek, W. H. Evers, B. Goris, A. D. Backer, A. P. Gantapara, F. T. Rabouw, J. Hilhorst, J. L. Peters, O. Konovalov, A. V. Petukhov, M. Dijkstra, L. D. Siebbeles, S. V. Aert, S. Bals, and D. Vanmaekelbergh, Nature Materials 15, 1248 (2016).
- Shevchenko et al. (2006) E. V. Shevchenko, D. V. Talapin, N. A. Kotov, S. O’Brien, and C. B. Murray, Nature 439, 55 (2006).
- Zhang et al. (2013) Y. Zhang, F. Lu, K. G. Yager, D. V. D. Lelie, and O. Gang, Nature Nanotechnology 8, 865 (2013).
- Jones et al. (2010) M. R. Jones, R. J. MacFarlane, B. Lee, J. Zhang, K. L. Young, A. J. Senesi, and C. A. Mirkin, Nature Materials 9, 913 (2010).
- Glotzer and Solomon (2007) S. C. Glotzer and M. J. Solomon, Nature Materials 6, 557 (2007).
- O’Brien et al. (2016) M. N. O’Brien, M. Girard, H. X. Lin, J. A. Millan, M. O. D. L. Cruz, B. Lee, and C. A. Mirkin, Proceedings of the National Academy of Sciences of the United States of America 113, 10485 (2016).
- Damasceno et al. (2012a) P. F. Damasceno, M. Engel, and S. C. Glotzer, Science 337, 453 (2012a).
- Gantapara et al. (2013) A. P. Gantapara, J. D. Graaf, R. V. Roij, and M. Dijkstra, Physical Review Letters 111, 1 (2013).
- Lee et al. (2023) S. Lee, T. Vo, and S. C. Glotzer, Nature Chemistry 15, 905 (2023).
- Karas et al. (2019) A. S. Karas, J. Dshemuchadse, G. V. Anders, and S. C. Glotzer, Soft Matter 15, 5380 (2019).
- Dullens and Petukhov (2007) R. P. Dullens and A. V. Petukhov, EPL 77 (2007), 10.1209/0295-5075/77/58003.
- Meijer et al. (2017) J. M. Meijer, A. Pal, S. Ouhajji, H. N. Lekkerkerker, A. P. Philipse, and A. V. Petukhov, Nature Communications 8 (2017), 10.1038/ncomms14352.
- Lu et al. (2021) J. Lu, X. Bu, X. Zhang, and B. Liu, Soft Matter 17, 6486 (2021).
- Keen and Goodwin (2015) D. A. Keen and A. L. Goodwin, Nature 521, 303 (2015).
- Chaney et al. (2021) D. Chaney, A. Castellano, A. Bosak, J. Bouchet, F. Bottin, B. Dorado, L. Paolasini, S. Rennie, C. Bell, R. Springell, and G. H. Lander, Physical Review Materials 5 (2021), 10.1103/PhysRevMaterials.5.035004.
- Meekel and Goodwin (2021) E. G. Meekel and A. L. Goodwin, CrystEngComm 23, 2915 (2021).
- Loidl (1989) A. Loidl, Annu. Rev. Phys. Chern 40, 29 (1989).
- Abbas et al. (2022) A. S. Abbas, E. Vargo, V. Jamali, P. Ercius, P. F. Pieters, R. M. Brinn, A. Ben-Moshe, M. G. Cho, T. Xu, and A. P. Alivisatos, ACS Nano 16, 9339 (2022).
- Vdovichenko et al. (2015) G. A. Vdovichenko, A. I. Krivchikov, O. A. Korolyuk, J. L. Tamarit, L. C. Pardo, M. Rovira-Esteva, F. J. Bermejo, M. Hassaine, and M. A. Ramos, Journal of Chemical Physics 143 (2015), 10.1063/1.4929530.
- Gantapara et al. (2015) A. P. Gantapara, J. D. Graaf, R. V. Roij, and M. Dijkstra, Journal of Chemical Physics 142 (2015), 10.1063/1.4906753.
- John et al. (2008) B. S. John, C. Juhlin, and F. A. Escobedo, Journal of Chemical Physics 128 (2008), 10.1063/1.2819091.
- Even et al. (2016) J. Even, M. Carignano, and C. Katan, Nanoscale 8, 6222 (2016).
- Beake et al. (2017) E. O. Beake, M. G. Tucker, M. T. Dove, and A. E. Phillips, ChemPhysChem 18, 459 (2017).
- Damasceno et al. (2012b) P. F. Damasceno, M. Engel, and S. C. Glotzer, ACS Nano 6, 609 (2012b).
- Deng et al. (2020) K. Deng, Z. Luo, L. Tan, and Z. Quan, Chemical Society Reviews 49, 6002 (2020).
- Elbert et al. (2021) K. C. Elbert, W. Zygmunt, T. Vo, C. M. Vara, D. J. Rosen, N. M. Krook, S. C. Glotzer, and C. B. Murray, Sci. Adv. 7, 23:abf9402 (2021).
- Simonov and Goodwin (2020) A. Simonov and A. L. Goodwin, Nature Reviews Chemistry 4, 657 (2020).
- Kundu et al. (2023) S. Kundu, K. Chakraborty, and A. Das, (2023).
- Onsager (1949) L. Onsager, Annals of the New York Academy of Sciences 51, 627 (1949).
- Alder and Wainwright (1957) B. J. Alder and T. E. Wainwright, Journal of Chemical Physics (U.S.) 27 (1957), 10.1063/1.1743957.
- Berryman (1983) J. G. Berryman, Physical Review A 27, 1053 (1983).
- Frenkel and Ladd (1984) D. Frenkel and A. J. Ladd, The Journal of Chemical Physics 81, 3188 (1984).
- Frenkel and Mulder (1985) D. Frenkel and B. M. Mulder, Molecular Physics 55, 1171 (1985).
- Veerman and Frenkel (1992) J. A. C. Veerman and D. Frenkel, Physical Review 45 (1992).
- McGrother et al. (1996) S. C. McGrother, D. C. Williamson, and G. Jackson, Journal of Chemical Physics 104, 6755 (1996).
- Bolhuis and Frenkel (1997) P. Bolhuis and D. Frenkel, Journal of Chemical Physics 106, 666 (1997).
- Haji-Akbari et al. (2009) A. Haji-Akbari, M. Engel, A. S. Keys, X. Zheng, R. G. Petschek, P. Palffy-Muhoray, and S. C. Glotzer, Nature 462, 773 (2009).
- Torquato and Jiao (2009) S. Torquato and Y. Jiao, Nature 460, 876 (2009).
- Agarwal and Escobedo (2011) U. Agarwal and F. A. Escobedo, Nature Materials 10, 230 (2011).
- Ni et al. (2012) R. Ni, A. P. Gantapara, J. D. Graaf, R. V. Roij, and M. Dijkstra, Soft Matter 8, 8826 (2012).
- Alloyeau et al. (2009) D. Alloyeau, C. Ricolleau, C. Mottet, T. Oikawa, C. Langlois, Y. L. Bouar, N. Braidy, and A. Loiseau, Nature Materials 8, 940 (2009).
- Gong et al. (2017) J. Gong, R. S. Newman, M. Engel, M. Zhao, F. Bian, S. C. Glotzer, and Z. Tang, Nature Communications 8 (2017), 10.1038/ncomms14038.
- Ong et al. (2017) W. L. Ong, E. S. O’Brien, P. S. Dougherty, D. W. Paley, C. F. Higgs, A. J. McGaughey, J. A. Malen, and X. Roy, Nature Materials 16, 83 (2017).
- Shen et al. (2019) W. Shen, J. Antonaglia, J. A. Anderson, M. Engel, G. van Anders, and S. C. Glotzer, Soft Matter 15, 2571 (2019).
- Anderson et al. (2016) J. A. Anderson, M. E. Irrgang, and S. C. Glotzer, Computer Physics Communications 204, 21 (2016).
V Supporting Information
V.1 Distribution of in the plastic phase
The distributions of calculated between the highest order rotational symmetry axes of the particles and all rotational symmetry axes of the crystal are shown in Figs. 7A,B,C for the three shapes, EPD, ESG and EPG. All the systems exhibit FCC crystals in the plastic phase where no orientational correlation is observed leading to the Gaussian distributions barring the statistical noise.
V.2 Orientation of the particles and relation with crystal symmetry
The fourteen particles of a random chosen FCC unit cell from the system of EPD, ESG and EPG are shown separately in fourteen sub-figures of Figs. 8, 9, 10 respectively; each sub-figure contains a single particle at the proper lattice site with similar color code for better visualization. Fig. 8 shows that the axes (the highest order rotational axis of symmetry) of each of the fourteen EPD particles are aligned in the directions of any eight axes of FCC crystal barring the statistical noise. The highest order symmetry axes () of ESG particles with point group also align with six axes of the FCC unit cells as per Fig. 9. Similarly, The axes of each EPG particle staying at the lattice site coincide with the direction of any eight axes of the crystal as shown in Fig. 10. The (hkl) planes are also shown in the unit cells which are responsible for the alignment of the particles. The color codes of all particles and axes were followed as defined in the main text.
V.3 Alignments of the lower order rotational symmetry axes of the particles with the axes of crystals
We explored the effect of lower order rotational symmetry axes of the particles in the alignment of the particles in the unit cells to get the insight about the orientational correlation present in the system. The EPD and ESG shapes contain the and axes respectively as the highest order rotational symmetry axes, but both the shapes have multiple axes present in the respective point groups as the lower order symmetry. The EPG shape contains only axes without having any other rotational symmetry operation in the point group. Figs. 11A,B shows the distributions of , and calculated between any of the axis of the particle and , and axes of the crystal separately for both shapes. The distributions suggest no direct relationship of the lower order rotational symmetry axes of the particle with the crystal structures where multiple crystal symmetry axes are involved without aligning in the directions of the particle axes.
V.4 Orientational ordered phase of Truncated Cuboctahedron (TC) system
Truncated Cuboctahedron (TC) shape was reported to self-assembly in body centred tetragonal (BCT) structure with tetragonal symmetry of crystallographic point group. In the densest state ( 0.74), the self-assembled TC system had two unique orientations with a fixed pairwise quaternion angle difference as shown in Fig. 12A. The TC shape contains six axes as the highest order rotational axes of point group and the tetragonal symmetry of the crystallographic point group consists of few rotational operations i.e., 2, , 2, 2. The two dimensional distribution of estimated between the six axes of the particle and any crystallographic rotational symmetry axes, shows the population density appears at with the axes of the tetragonal crystal. Other particles in the system are found to create angles and with the crystal axes as shown in Fig. 12B, where all the , , axes are considered as equivalent. The existence of two unique orientations in the system leaves the signatures of population densities in the histogram at two different values as confirmed by the Fig. 12B. This indicates that all the highest order particle symmetry axes are not required to align in the directions of any crystallographic rotational axes to maintain the orientationally ordered crystalline phase as confirmed by the appearance of the peaks at shown in Fig. 12B.